# Unsupervised Episode Detection for Large-Scale News Events

### v1

@@ 5. Experimental Setup
Table 4. Compares the top-5 salient terms which (1) have the highest cosine-similarity (CS) and (2) highest discriminative co-occurrence (DC), with the given keyword.
@@
Table 5. Gold and detected episodes (a maximum of five are included for brevity) for the “2019 Hong Kong Legislative Protests” key event. We specify the gold/detected episode attributes for each episode cluster in the following semicolon-separated format: core entity; action; object; time; location. “Not detected” denotes that no more episodes were generated by the model.
@@
Hong Kong protesters; storm Legislative Council; government and police; July 1, 2019; Legislative Council Complex, Hong Kong Hong Kong citizens; march against extradition bill; ”Carrie Lam and Chinese government; June 2019; Various locations in Hong Kong Not detected Not detected
@@
EpiMine (total 7 eps) protesters; broke into and occupied; Hong Kong’s legislative building; July 1, 2019; Hong Kong protesters; vandalized; the legislative building; after breaking in; Hong Kong police; fired tear gas at; protesters; after midnight on July 1; outside the legislative building Carrie Lam; condemned; the protesters’ actions; in a news conference at 4am on July 2; Hong Kong police; began making arrests of; protesters involved; in the days after; Hong Kong
5.1. Datasets
We conduct our experiments on three thematic, real-world news corpora selected from Wikipedia [^3] https://en.wikipedia.org/wiki/ over the last decade. For each theme, we manually collect approximately $10$ key events composed of multiple articles and ensure that distinct episodes exist in these articles. These articles are originally obtained from the hyperlinks of references from the Wikipage of each key event, and then filtered with constraints in time, language, and relevance. We avoid excessive filtering (allowing for articles that may contain either a subset or even none of the ground truth episodes identified) in order to preserve the challenging, real-world nature of the task.
During the collection process, we targeted selecting a diverse set of key events topics within a theme. For instance, we attempted to cover every type of “natural disaster”, including tornados, wildfires, and etc. When selecting key events, we leave out those with less than $20$ hyperlinks in the Wikipage. Table 1 summarizes the statistics for these datasets. We also construct a background news corpus of approximately 4,000 long news articles using the New York Times corpus for topic categorization (Meng et al., 2020). We provide further details on the dataset collection criteria/process, each theme, and corresponding key events in Appendices E and F.
5.2. Baseline Models
We compare against the following methods using the evaluation metrics specified in Appendix D. For the two most competitive baselines, K-means and EvMine, we also integrate EpiMine’s same exact LLM model/prompt (Appendix C) for candidate episode estimation (Section 4.3) and confidence-based episode-segment classification process (Section 4.4) to demonstrate any LLM-based improvements.
@@ 5.2. Baseline Models
  • K-means (Likas et al., 2003): No. of ground-truth episodes is given; clusters segments based on semantic similarity of ST  (Reimers and Gurevych, 2019) embeddings.
@@
  • EvMine (Zhang et al., 2022): Unsupervised framework for key event detection that leverages peak phrases and detects communities using event-indicative features. We extend the original document-level method to the segment level for episode detection.
@@
  • EMiner (Jiao et al., 2023): Unsupervised event chain mining that performs atomic action clustering. For episode detection, we map its final output, a list of events, back to the original sentences from which each event was extracted, treating these sentences as segments. To retrieve more episode-associated segments, we use ST (Reimers and Gurevych, 2019) to select the $k$ most similar segments to each cluster sentence.
@@
5.3. Overall Results & Analysis
In Table 2, EpiMine shows an average 80.8% increase in 5-precision, a 34.0% increase in 5-recall, and a 62.8% increase in 5-F1 over all baselines. Notably, despite both ’K-means’ and ’K-means + LLM’ being given the ground-truth number of episodes, they are significantly outperformed by EpiMine (both the base model and no confidence ablation). Additionally, EvMine and EMiner, originally designed for key event and atomic action levels of event granularity, fail to address the unique challenges of episode detection. We further analyze our results through extensive quantitative and qualitative studies, including a detailed case study on the “2019 Hong Kong Legislative Protest” (as shown in Figure 1), leading to the following takeaways:
  1. (1)
    The strengths of discriminative co-occurrence complement those of cosine similarity:
    Table 4 illustrates the qualitative strengths of our novel discriminative co-occurrence metric. Both cosine similarity (CS) and discriminative co-occurrence (DC) offer different, complementary strengths. CS identifies similar words that play a similar role or are synonyms within an episode (e.g., “broke”, “ransacked”), while DC identifies the key surrounding actions and objects that co-occur within the same episode (e.g., “slogans”, “wall”). This is quantitatively supported by the ablation studies in Table 3, which show a significant decrease in the quality of our top articles (partitioned into episodes) without our salience and discriminative co-occurrence measures. The politics dataset does show slight improvements in precision and F1 when discriminative co-occurrence is replaced, due to more term overlap across episodes, resulting in less distinct, sequential episodes.
  2. (2)
    LLMs require effective episode fragment clusters for quality episode estimation:
    In Table 5, we compare EpiMine with the most competitive baselines integrated with our LLM-based candidate episode estimation and classification, as well as Claude-2.1 and ChatGPT-$4$https://chat.openai.com/. Without any initial clusters as guidance, LLMs alone (Claude, GPT-4) fail to detect high-quality episodes, missing most ground-truth episodes and including irrelevant atomic actions (e.g., “Brian Leung pulls off mask”). Similarly, using low-quality clusters from the baselines results in poor performance. EvMine detects episodes that all reflect the same event, “Protesters vandalized the Legislative Chamber”. While K-means produces more distinct episodes, it does not capture the most critical, gold episodes. In contrast, EpiMine’s episodes are both distinct and meaningful.
    An interesting observation is that EpiMine’s clusters elicit the LLM to identify more meaningful temporal information. Unlike most baseline episodes, which simply have “July 1, 2019” as the time attribute, EpiMine’s episodes feature more descriptive temporal cues: “after breaking in”, “after midnight”, “in a news conference at 4 am on July 2”. Moreover, EpiMine’s “incorrect” episode #4 is a significant sub-event of the key event discussed by many articles. This strongly demonstrates the impact of EpiMine’s candidate episode clusters as input into the LLM; LLMs alone cannot perform quality episode detection.
  3. (3)
    Fragment ranking effectively identifies top articles:
    In Figure 4, we conduct a sensitivity analysis of the top articles chosen to estimate candidate episodes (Section 4.3). We compare the gold episodes contained in the set of the top $\delta$articles as we vary $\delta$. By ranking the articles based on their likelihood of containing both high-quality and numerous episodes, we find that EpiMine’s top article selection covers the vast majority of episodes by $\delta=25\%$ and more comprehensively around $\delta=45\%$. This is significant as it helps minimize both the noise and the amount of data needed to accurately detect all episodes.
  4. (4)
    Confidence leads to more conservative classification:
    The “No-Confidence” ablation in Table 2 shows that using confidence (Equation 3) improves EpiMine’s precision but reduces recall. This conservative approach results in EpiMine being cautious when assigning segments to episode clusters, excluding segments with insufficient confidence. This indicates that our method relies on both our clustering method and confidence scoring for precise episode detection. EpiMine users can determine if they prefer a confidence or no-confidence method depending on their use-case (higher precision versus higher recall). Nonetheless, both version still outperform all baselines.
Figure 4. Percentage of key event’s gold episodes captured in the $\delta$top articles chosen during the candidate episode estimation. Results averaged across themes.
6. Conclusion
In this work, we propose our novel method EpiMine for unsupervised episode detection for large-scale news events. EpiMine performs (1) episode indicative term mining, which identifies combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes, (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters. EpiMine significantly outperforms all baselines on the vast majority of key events, as shown through extensive quantitative and qualitative analysis. Further work towards the temporal analysis of episodes within articles can be explored, as well as extending our work to primarily multilingual news settings with low resources.
Appendix A Ethics Statement
Based on our current methodology and results, we do not expect any significant ethical concerns, given that subtasks like episode detection within the news event extraction and analysis is a standard problem domain across data mining applications. Furthermore, having the method rely on zeor supervision helps as a barrier to any user-inputted biases. However, one minor factor to take into account is any hidden biases that exist within the large language models used as a result of any potentially biased data that they were trained on. We used these pre-trained language models for refining the fluency of the detected episode clusters and did not observe any concerning results, as it is a low-risk consideration for the domains that we studied.
Appendix B Additional Details for Episode Partitioning
@@
Finally, following (Wang et al., 2021; Kargupta et al., 2023), we take the harmonic mean of all pairwise discriminative co-occurrence scores instead of a simple average. This allows us to prioritize the more salient and discriminative terms when determining the episode partitions. For instance, if “protesters” consistently occurs throughout the majority of episodes and thus has a low average discriminative co-occurrence, then it is not as informative for episode partitioning.
Appendix C Claude-2 Prompt & Example for Candidate Episode Estimation
@@
Appendix D Evaluation Metrics
Following a recent work on key event detection (Zhang et al., 2022), we adapt the $k$-prec, $k$-recall, and $k$-F1 to quantitatively evaluate the episode detection results. We use these metrics to evaluate how the model output matches the ground truth episodes using the top-$k$ segments within each detected episode. Formally, suppose there are $N$ ground truth episodes $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$, each of which is a set of text segments related to its corresponding episode. =$\mathcal{E}=\{E_{1},E_{2},\dots,E_{K}\}$ are the model predicted episodes, each of which is a ranked list of segments, and $E_{j,k}$ means the top-$k$ segments within $E_{j}$. Then, the $k$-metrics are defined as follows:
k-prec $=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{\sum_{E_{j}\in\mathcal{E}}\mathds{1}(|E_{j}|\geq k)}$
k-recall $=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{N}$
k-F1 $=\frac{2\cdot\text{k-prec}\cdot\text{k-recall}}{\text{k-prec}+\text{k-recall}}$
Appendix E Key Event Corpus Retrieval
Given that our task is novel and no large-scale key event-specific news corpus is available for this task where the key events are guaranteed to contain distinguishable episodes, we briefly discuss how we collect the input corpus from online news data, even if it is out of scope for our fundamental methodology. Given our set of key events (as listed in Section 5.1), we first scrape the external reference list from their corresponding Wikipedia page and select the news articles that have been published within two months of given key event’s start date (e.g., all articles selected for “January 6 2021 Capitol Attack” would have been published between November 6-March 6). This is important as we want to prioritize the news articles which focus on describing the episodes of the key event and their corresponding aspects as opposed to primarily opinions or analyses. This allows us to motivate our task as one critical for currently evolving key events which required a more fine-grained episodic timeline. Furthermore, it is consequently unlikely for a single article to cover all of the episodes and exclusively episodes under a key event. Despite this being more challenging, it is acceptable as the goal of our task is to extract only the key event-related episodes, which must be substantiated by multiple documents in either case. We list all of our selective themes and their corresponding key events included in the dataset that we construct:
  • Natural disasters: 2023 Tornado outbreak sequence; 2023 Hawaii Wildfires; 2021 Western Kentucky tornado; 2017 Mocoa landslide; 2010 Haiti earthquake; 2021 Henan floods; 2019 Nyonoksa radiation accident; 2022 North American winter storm; 2011 Fukushima nuclear accident
  • Political Events: 2020 Kyrgyz Revolution, 2019 Storming of the Hong Kong Legislative Council Complex; 2019 Siege of the Hong Kong Polytechnic University; 2017 Zimbabwean coup; 2018 Italian government formation; 2021 January 6 United States Capitol attack; 2018 Thai Cave Rescue Operation; 2018 Armenian Revolution; 2017 Lebanon–Saudi Arabia dispute; 2013 Tunisian political crisis
Appendix F Claude-2 Prompt for Dataset Annotation

### v2

@@ 5.2 Baselines
Table 4: Compares top-5 salient terms which (1) have the highest cosine-sim (CS) and (2) discriminative co-occurrence (DC), with the given keyword.
@@
Table 5: Gold and detected episodes (a maximum of five are included for brevity) for the “2019 Hong Kong Legislative Protests” key event. We specify the gold/detected episode attributes for each episode cluster in the following semicolon-separated format: core entity; action; object; time; location. “Not detected” denotes that no more episodes were generated by the model. We note the number of detected episodes beside the model name. We also color-code attributes which clearly align to a specific episode.
@@
Hong Kong protesters; storm Legislative Council; government and police; July 1, 2019; Legislative Council Complex, Hong Kong Hong Kong citizens; march against extradition bill; "Carrie Lam and Chinese government; June 2019; Various locations in Hong Kong Not detected Not detected
EpiMine w/ [Uncaptioned image]
(7 eps)
@@
Protesters; broke into and occupied; Hong Kong’s legislative building; July 1, 2019; Hong Kong Protesters; vandalized; the legislative building; after breaking in; Hong Kong Police; fired tear gas at; protesters; after midnight on July 1; outside the legislative building Carrie Lam; condemned; the protesters’ actions; in a news conference at 4am on July 2; Hong Kong Police; began making arrests of; protesters involved; in the days after; Hong Kong
5.3 Overall Results & Analysis
In Table 2, EpiMine shows an average 80.8% increase in 5-precision, a 34.0% increase in 5-recall, and a 62.8% increase in 5-F1 over all baselines. Notably, despite both K-means and K-means + [Uncaptioned image] being given the ground-truth number of episodes, they are significantly outperformed by EpiMine (both the base model and no confidence ablation). Additionally, EvMine and EMiner, originally designed for key event and atomic action levels of event granularity, fail to address the unique challenges of episode detection. We further analyze our results through extensive quantitative and qualitative studies, including a detailed case study on the “2019 Hong Kong Legislative Protest” (as shown in Figure 1), leading to the following takeaways:
1. LLMs require effective episode fragment clusters for synergistic episode estimation. As shown in Table 5, LLMs without any initial clusters as guidance ([Uncaptioned image] , GPT-4333https://chat.openai.com/) fail to detect high-quality episodes, miss most ground-truth episodes, and include irrelevant atomic actions (e.g., “Brian Leung pulls off mask”). Similarly, using low-quality baseline clusters results in poor performance. EvMine detects episodes that all reflect the same event, “Protesters vandalized the Legislative Chamber”. While K-means produces more distinct episodes, it does not capture the most critical, gold episodes. In contrast, EpiMine’s episodes are both distinct and meaningful, attributed to its cluster quality. This is quantitatively confirmed by EpiMine-No LLM’s competitive performance: using only EpiMine’s episode fragment clusters to compute episode representations—without any LLM summarization—still yields significantly better performance than all baselines (without or without LLM integration) on the Terrorism and Politics datasets, and remains highly competitive on Natural Disasters. This indicates the high quality of our fragment ranking and clustering.
Figure 4: Percentage of key event’s gold episodes captured in the $\delta$top articles chosen during the candidate episode estimation. Results averaged across themes.
EpiMine’s clusters also elicit the LLM to identify more meaningful temporal information. Unlike most baseline episodes which have “July 1, 2019” as the time attribute, EpiMine’s episodes feature more descriptive temporal cues: “after breaking in”, “after midnight”, “in a news conference at 4 am on July 2”. Moreover, EpiMine’s “incorrect” episode #4 is a significant sub-event of the key event discussed by many articles. This strongly demonstrates the impact of EpiMine’s candidate episode clusters as input into the LLM; LLMs alone cannot perform quality episode detection.
2. The strengths of discriminative co-occurrence complement those of cosine similarity. Table 4 illustrates the qualitative strengths of our novel discriminative co-occurrence metric. Both cosine similarity (CS) and discriminative co-occurrence (DC) offer different, complementary strengths. CS identifies similar words that play a similar role or are synonyms within an episode (e.g., “broke”, “ransacked”), while DC identifies the key surrounding actions and objects that co-occur within the same episode (e.g., “slogans”, “wall”). This is quantitatively supported by TF-IDF and No DC in Table 3, which show a significant decrease in the quality of our top episode clusters without our salience and discriminative co-occurrence measures. We note that the politics dataset does show slight improvements in precision and F1 when discriminative co-occurrence is replaced, likely due to more term overlap across episodes (less distinct, sequential episodes).
3. Fragment ranking identifies top articles. In Figure 4, we conduct a sensitivity analysis of the top articles chosen to estimate candidate episodes (Section 4.3). We compare the gold episodes contained in the set of the top $\delta$articles as we vary $\delta$. By ranking the articles based on their likelihood of containing both high-quality and numerous episodes, we find that EpiMine’s top article selection covers the vast majority of episodes by $\delta=25\%$ and more comprehensively around $\delta=45\%$. This is significant as it helps minimize both the noise and the amount of data needed to accurately detect all episodes. This is further supported by the EpiMine-Top ablation (Table 3), which quantitatively shows that the top 25% of the ranked fragments alone have competitive or higher performance than the baselines.
The Role of Confidence. The confidence metric influences EpiMine toward more conservative episode-segment classification by pruning segments with statistically insignificant confidence scores ($\leq\mu-\sigma$), which are unlikely to map to a single episode. Reducing this threshold increases the number of mapped segments, potentially improving recall by converting false negatives into true positives, but at the cost of reduced precision due to an increase in false positives. This trade-off is quantitatively demonstrated in the “No-Confidence” ablation (Table 2), where omitting confidence leads to occasional gains in recall accompanied by declines in precision. Nevertheless, both configurations—with and without confidence—are significantly better than all baselines, allowing users to determine if including confidence aligns with their use case.
6 Conclusion
In this work, we proposed EpiMine, a novel, unsupervised episode detection method for large-scale news events. EpiMine performs (1) episode indicative term mining— identifying combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes, (2) episode partitioning, which partitions each article into approximate isolated episodes, (3) LLM-enhanced episode estimation, which clusters the top partitions into candidate episodes and synergizes with an LLM to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters. EpiMine significantly outperforms all baselines on the vast majority of key events, as shown through extensive quantitative and qualitative analysis.
7 Limitations & Future Work
While EpiMine serves as an intuitive, unsupervised framework which demonstrates a more interpretable granularity for event analysis (episodes), it contains a few limitations that form the foundation for future, impactful research areas.
We note that the key event theme has an impact on EpiMine’s performance. Specifically, natural disaster episodes are typically sequential and semantically distinct: disaster begins $\rightarrow$ warning $\rightarrow$ evacuation $\rightarrow$ damage/deaths $\rightarrow$ relief. As K-means is uniquely given $k$, the number of episodes, and relies on semantic similarity, it performs well with distinct episodes. However, we still see that its reliance on surface-level semantics leads to lower precision. Additionally, in the ablation studies shown in Table 3, the politics dataset does show slight improvements in precision and F1 when discriminative co-occurrence is replaced, due to more term overlap across episodes, resulting in less distinct, sequential episodes.
Further work towards the temporal analysis of episodes within articles can be explored, as well as extending our work to primarily multilingual news settings with low resources.
8 Ethics Statement
Based on our current methodology and results, we do not expect any significant ethical concerns, given that subtasks like episode detection within the news event extraction and analysis is a standard problem domain across data mining applications. Furthermore, having the method rely on zero supervision helps as a barrier to any user-inputted biases. However, one minor factor to take into account is any hidden biases that exist within the large language models used as a result of any potentially biased data that they were trained on. We used these pre-trained language models for refining the fluency of the detected episode clusters and did not observe any concerning results, as it is a low-risk consideration for the domains that we studied.
9 Acknowledgements
This work was supported by the National Science Foundation Graduate Research Fellowship. This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois. DeltaAI is a joint effort of the University of Illinois at Urbana-Champaign and its National Center for Supercomputing Applications.
Appendix A Experimental Settings
For implementing EpiMine, we use the following hyperparameters across all datasets: $\delta=25\%$, $sim\_thresh=0.75$. All other hyperparameters are set to their respective default values. We provide all experimental settings in Appendix A.To determine statistical significance, we check for $\geq\mu-\sigma$. For our word representations, we use bert-base-uncased. For our sentence representations, we use all-mpnet-base-v2. We choose to use Claude-2.1444claude.ai/ for fluent candidate episode estimation due to its strong structured JSON/XML input and output formatting abilities. However, this proprietary model can be replaced with any open-source model as EpiMine is model-agnostic. We use only one NVIDIA GeForce GTX 1080 for all experiments; for non-API models, we utilize two NVIDIA-RTX A6000s.
Appendix B Baselines
We compare against the following methods using the evaluation metrics specified in Appendix C.
@@ Appendix B Baselines
  • K-means Likas et al. (2003): No. of ground-truth episodes is given; clusters segments based on semantic similarity of ST  Reimers and Gurevych (2019) embeddings.
@@
  • EvMine Zhang et al. (2022): Unsupervised framework for key event detection that leverages peak phrases and detects communities using event-indicative features. We extend the original document-level method to the segment level for episode detection.
@@
  • EMiner Jiao et al. (2023): Unsupervised event chain mining that performs atomic action clustering. For episode detection, we map its final output, a list of events, back to the original sentences from which each event was extracted, treating these sentences as segments. To retrieve more episode-associated segments, we use ST Reimers and Gurevych (2019) to select the $k$ most similar segments to each cluster sentence.
@@
  • No LLM: We take the estimated episode clusters from Section 4.3 that normally would have been inputted into the LLM, and instead use them to compute our episode representations directly. These representations are used for our classification step (Section 4.4), run on the full dataset with confidence.
@@
Appendix C Evaluation Metrics
Following a recent work on key event detection Zhang et al. (2022), we adapt the $k$-prec, $k$-recall, and $k$-F1 metrics to quantitatively evaluate episode detection performance— specifically, how the model’s top-$k$ segments within each detected episode align with the ground truth episodes.
Formally, suppose there are $N$ ground truth episodes $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$, each of which is a set of text segments related to its corresponding episode. =$\mathcal{E}=\{E_{1},E_{2},\dots,E_{K}\}$ are the model predicted episodes, each of which is a ranked list of segments, and $E_{j,k}$ means the top-$k$ segments within $E_{j}$. Then, the $k$-metrics are defined as follows:
$\begin{split}\text{{k-prec}}&=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{\sum_{E_{j}\in\mathcal{E}}\mathds{1}(|E_{j}|\geq k)}\\ \text{{k-recall}}&=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{N}\\ \text{{k-F1}}&=\frac{2\cdot\text{k-prec}\cdot\text{k-recall}}{\text{k-prec}+\text{k-recall}}\\ \end{split}$
Appendix D Key Event Corpus Dataset Construction & Annotation
Given that our task is novel and no large-scale key event-specific news corpus is available for this task where the key events are guaranteed to contain distinguishable episodes, we briefly discuss how we collect the input corpus from online news data. Given our set of key events (as listed in Section 5.1), we first scrape the external reference list from their corresponding Wikipedia page and select the news articles that have been published within two months of given key event’s start date (e.g., all articles selected for “January 6 2021 Capitol Attack” would have been published between November 6-March 6). This is important as we want to prioritize the news articles which focus on describing the episodes of the key event and their corresponding aspects as opposed to primarily opinions or analyses. This allows us to motivate our task as one critical for currently evolving key events which required a more fine-grained episodic timeline. Furthermore, it is consequently unlikely for a single article to cover all the episodes and exclusively episodes under a key event. Despite this being more challenging, it is acceptable as the goal of our task is to extract only the key event-related episodes, which must be substantiated by multiple documents in either case.
During the collection process, we targeted selecting a diverse set of key events topics within a theme. For instance, we attempted to cover every type of “natural disaster”, including tornados, wildfires, and etc. When selecting key events, we leave out those with less than $20$ hyperlinks in the Wikipage and manually inspect at least 20 articles per event in order to ensure quality. Table 1 summarizes the statistics for these datasets. We also construct a background news corpus of approximately 4,000 long news articles using the New York Times corpus for topic categorization Meng et al. (2020).
For the annotation process, we had each of the four individual annotators (computer science graduate students) manually identify a ground-truth description for each of the episodes under every key event. The descriptions and a one-shot annotation demonstration (assigning either an episode ID per segment, or ‘M’/‘X’ if it describes multiple or none) are provided in our automatic annotation prompt (Appendix I).
The same annotators each manually annotated a subset of segments (25 articles, 300 segments in total) with either their corresponding episode ID or ‘-1’ if the segment was either ‘X’ and ‘M’. The human-automatic annotation agreement is shown in Table 6 with three versions of intra-class correlation (ICC) and Cohen’s $\kappa$). The Cohen’s $\kappa$ indicates substantial agreement, and the all three versions of ICC indicate good reliability.
Cohen’s $\kappa$ ICC1k ICC2k ICC3k
Score 0.614 0.772 0.772 0.773
Table 6: Agreement scores between the human and LLM annotation of each article segment.
Appendix E Identifying Salient Terms for Episode Detection
$\begin{split}\textbf{Salience}(w_{i})=&\left(1+\log^{2}\left(freq(w_{i})\right)\right)\\ &\times\log\left(\frac{N_{bg}}{bgf(w_{i})}\right)\end{split}$ (3)
Appendix F Key Event News Article Pre-Processing
Given that the expected output for the episode detection task is a cluster of text segments, we first must segment each key event news article. We would like to ideally preserve both the primary aspects (e.g., core entities and their actions) and peripheral aspects (e.g., reactions to a core entity’s action) relevant to that episode, which may be helpful for cross-document episode co-reference resolution. In order to do this, we utilize the text segmentation method, C99 Choi (2000). Furthermore, in order to assist with the cohesiveness of the segment, we employ entity co-reference resolution before performing segmentation, which assists with retaining the context across text segments (“They surrounded the legislative building […]” $\rightarrow$ “The protesters surrounded the legislative building […]”). Our core methodology is given these text segments (in their raw form, without co-references resolved) and their source articles as the primary inputs.
Appendix G Additional Details for Episode Partitioning
@@
Finally, following Wang et al. (2021); Kargupta et al. (2023), we take the harmonic mean of all pairwise discriminative co-occurrence scores instead of a simple average. This allows us to prioritize the more salient and discriminative terms when determining the episode partitions. For instance, if “protesters” consistently occurs throughout the majority of episodes and thus has a low average discriminative co-occurrence, then it is not as informative for episode partitioning.
Appendix H Claude-2 Prompt & Example for Candidate Episode Estimation
@@
Appendix I Claude-2 Prompt for Dataset Annotation